The flux of magnetic field through a closed conducting loop changes with time according to the equation, Φ = at^2 + bt + c. (a) Write the SI units of a, b and c. (b) If the magnitudes of a, b and c are 0.20, 0.40 and 0.60 respectively, find the induced emf at t = 2 s.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. ϕ = at2 + bt + c (a) a = [ϕ/T^2] = [ϕ/t/t] = Volt /Sec b = [ϕ/t] = Volt c = [ϕ] = Weber (b) E = d ϕ/dt [a= 0.2, B = 0.4 , 0.6,, t = 2s] = 2at + b